### v1

# Heliometric stereo: a new frontier in surface profilometry

Abstract
Accurate and reliable measurements of three-dimensional surface structures are important for a broad range of technological and research applications, including materials science, nanotechnology, and biomedical research. Scanning helium microscopy (SHeM) uses low-energy () neutral helium atoms as the imaging probe particles, providing a highly sensitive and delicate approach to measuring surface topography. To date, topographic SHeM measurements have been largely qualitative, but with the advent of the heliometric stereo method –- a technique that combines multiple images to create a 3D representation of a surface –- quantitative maps of surface topography may now be acquired with SHeM. Here, we present and discuss two different implementations of heliometric stereo on two separate instruments, a single detector SHeM and a multiple-detector SHeM. Both implementations show good accuracy (5% and 10% respectively) for recovering the shape of a surface. Additionally, we discuss where heliometric stereo is most applicable, identify contrast features that can limit its accuracy, and discuss how to mitigate these limitations with careful design and sample choices that be readily implemented on current instruments.
I Introduction
Accurate measurements of surface topography (often termed surface metrology) are crucial to modern research and development, with widespread applications from understanding material behavior to optimizing manufacturing[1, 2]. Significantly, surface metrology can also provide assurance that products and materials meet required standards for functionality, quality, and safety[3]. As new technologies emerge, for example, in micro and nanoscale organic and quantum devices, reliable measurement of surface topography on an ever smaller scale is crucial. Scanning helium microscopy (SHeM)[4, 5, 6, 7, 8] is a technique that uses a narrow beam of neutral helium atoms (), generated via a collimating pinhole to produce topographic micrographs. The use of a neutral atom beam has two key advantages for topographic measurements. First, it delivers an incident low energy beam without the resolution being diffraction limited: the typical beams used in SHeM offer a de Broglie energy of , corresponding to a wavelength of , [9, 10], rendering beam damage, as observed with charged beam methods[11], impossible. Second, thermal helium atoms scatter off the outermost electron density distribution, located above the ion cores of surface atoms, which results in information exclusively from the surface, without any contribution due to beam penetration into the bulk[9]. Existing scattering techniques for performing surface metrology, such as optical profilometry, laser scanning confocal microscopy, scanning electron microscopy, interferometric profilometry and structured light scanning[12, 13, 14], either use highly energetic probe particles, which can induce changes in the sample, or can have a non-negligible interaction volume with the surface and thus do not measure the surface topography exclusively. Contact techniques, such as stylus profilometers, scanning tunneling microscopes and atomic force microscopes[15, 16, 17, 18], provide a more direct surface profile measurement through contact with a physical probe. However, these are largely limited in either lateral resolution (stylus profilometers) or in aspect ratio (AFM/STM), can induce changes in the sample due to the direct contact with the probe, and in all cases convolution with the probe tip must be considered.
SHeM has demonstrated its ability to measure surface topography qualitatively[19, 20, 21, 22], and two previous publications have shown SHeM to be capable of providing individual pieces of 3D information. Myles et al.[23] applied stereophotogrammetry to taxonomy, measuring specific dimensions of biological specimens. Lambrick et al.[24] demonstrated that quantitative 3D measurements can be made by interpreting specific contrast features of SHeM to depth profile trenches in silicon precisely. Both approaches make accurate measurements of specific dimensions, but do not yield a 3D view of the entire sample.
In this paper, we describe the development of 3D imaging using SHeM (a technique termed heliometric stereo) using both single detector [25], and multiple detector instruments. We also provide detailed analysis of heliometric stereo reconstruction accuracy in general, how it is affected by contrast mechanisms unique to SHeM, and how these sources of error can be mitigated using instrumentation, algorithms and/or careful sample selection[26]. Sections II, III and IV restate the heliometric stereo method, supporting theory and surface reconstructions using a single detector, respectively, as presented by Radić et al.[25] with more detail than originally presented.
II Heliometric Stereo
Analogously to photometric stereo in optics, heliometric stereo is based on the assumption that a given point on a surface will scatter the incident helium beam with a known scattering distribution, formally termed the bidirectional reflection distribution function (BRDF)[27]. This distribution defines the scattered intensity as a function of the incoming and outgoing angles of the helium beam relative to the surface. If the microscope detector is fixed at a known angle to the surface and the incident direction is known, then the recorded intensity is solely a function of surface orientation and scattering distribution. SHeM instruments have a well-defined scattering geometry[28], which gives a direct relationship between the intensity of each pixel in a micrograph and the local surface orientation of the corresponding point on the sample.
Recent experimental results have shown that the majority of unprepared surfaces studied in SHeM exhibit almost exclusively diffuse, or close to diffuse, scattering [28, 24, 20, 19, 29, 30] enabling, crucially, the first experimental implementation of heliometric stereo[25]. These materials have typically been unprepared ‘technological samples’, in contrast to conventional surface science studies where surfaces are prepared and maintained in a pristine, atomically perfect state, and where specular or Bragg scattering is often dominant[9, 10, 31, 32, 33, 34]. The observation of diffuse scattering from technological surfaces arises from the fact that the de Broglie wavelength of the helium atoms () is comparable to the interatomic spacing in solids. As such, given that most technological samples have atomically rough surfaces, diffuse scattering dominates. Consequently, the observed diffuse distribution is consistent with Knudsen’s cosine law[35], and is analogous to Lambert’s cosine law for visible light[36].
III Theory
The following derivation is a summary of the heliometric stereo method previously presented by Lambrick & Palau et al.[26] Mathematically, diffuse scattering follows a cosine distribution centered on the surface normal,
where is the angle between the local surface normal, , and the outgoing scattering direction . It can be shown that for cosine scattering with a macroscopic circular detection area – such as is present experimentally in SHeM[5, 37] – that equation (1) holds the same mathematical form up to a multiplicative constant[26]. Equation (1) has the same form for light in photometric stereo[38], however, due to differences in image projection for light, represents the illumination direction (the outgoing scattering direction). The equivalence between detection and illumination in SHeM and in ordinary optical imaging are shown in figure 1. With light, there is a broad illumination from a light source and the scattered light is focused to form an image in the camera. In SHeM a focused beam is produced that scatters from the sample with a selection of the scattered atoms detected, the focused beam is then scanned over the sample to create an image via Helmholtz reciprocity. With light, Lambertian cosine scattering[36] places a cosine dependence on the angle between the illumination direction and the surface normal, which leads to the same equation (1), just with a different definition of . Both heliometric- and photometric-stereo then follow the same mathematical procedure to produce a topographic reconstruction.
As equation (1) directly relates the intensity scattered in a particular direction to the orientation of the surface, it is possible to acquire the surface orientation from a series of measured intensities. In SHeM that means multiple different detection directions, for the same illumination direction, see an example in figure 1; for photometric stereo with light that would mean multiple separate illumination directions with the same camera position.
For multiple detection directions in a SHeM system, equation (1) may be written in matrix form,
where is a -dimensional vector of pixel intensities, all for the same position on the sample, corresponding to images taken with different detection directions. is a matrix containing the normalized vectors pointing from the scattering point to the detectors. The linear system is thus solved,
Assuming the height of the surface can be described by a continuous function of the lateral position, i.e. , then
Thus once the surface normals are found, the gradient field given by equation (5) may be integrated to obtain an equation of the surface profile, i.e. a topographic map of the sample. A regularized least squares approach, developed by Harker and O’leary[39, 40], is used in the current work.
The orientation of a surface can be defined by two angles, plus, following the formalism of photometric stereo, we allow for a constant of proportionality, the albedo factor ( in equation (1)), hence there are 3 degrees of freedom. Therefore, a minimum of 3 data points (detector directions) are needed for each position on the surface to reconstruct the 3D topography. In practice SHeM data is often signal-to-noise limited and there may be minor deviations to the cosine law, thus at least 4 are recommended, but as demonstrated by our results in section V, not necessarily required, for a reconstruction to be robust. Indeed, the method has proved robust to noise in simulated SHeM micrographs[26].
IV Reconstruction using a single-detector instrument
A key result found in previous work[26] was that rotations about the beam axis in SHeM may be used to create multiple ‘effective detection directions’, on an instrument with a single detector. In figure 2 that acquisition process is illustrated: an image is taken with the sample in one orientation (a), then the sample is rotated and a second image is taken (b). As the rotation occurred around the beam axis, the rotation of the sample is equivalent to a rotation of the detector (c). Therefore multiple detection directions may be sampled in an instrument with a single physical detector (d). However, if the rotation is performed about an axis other than one parallel to the beam axis the illumination of the sample is changed, and thus the correspondence between sample rotations and multiple detectors is lost [26].
The single detector SHeM at Cambridge (A-SHeM), was designed to operate with an incidence angle of [5] enabling topographic imaging as well as studies of specular scattering. For the current work we modified the A-SHeM to operate in a normal incidence configuration through a new 3D printed ‘pinhole plate’ optical element[41, 42]. The pinhole plate mounts the collimating pinhole, defines the detection geometry and, in the A-SHeM, is used to mount the sample and sample manipulation stages. The old () (a) and new () (b) configurations are shown in figure 3, with the incident helium beam and rotation axis highlighted. The new pinhole plate macroscopically changes the mounted orientation of both the sample and the sample manipulation stages relative to the incident beam: in figure 3 (b) the beam is now incident at to the sample and the axis of rotation is parallel to the beam. However, while the rotation axis and the beam are parallel, they are not coaxial, thus an alignment procedure was needed to ensure that the reconstructed image area remained centered in the micrographs as the sample was rotated. The same rotation and alignment procedure as used by von Jeinsen et al.[10] was applied here. While the the new design does introduce a small () change in the path length from the collection aperture to the detector (figure 3), the detection efficiency, and therefore image acquisition, is unaffected[20].
Panel 1 adapted from a figure by T. Piesk, CC BY 4.0, published 2018, https://commons.wikimedia.org/w/index.php?curid=66347083.
To test the reconstruction accuracy for the A-SHeM implementation, a sample of well-defined topography was needed that matched the basic assumptions of heliometric stereo, principally a continuous surface. Aluminum potassium sulfate crystals () were chosen. They possess a well-defined octahedral structure and can be readily grown to form single crystals on micron to millimeter length-scales. To demonstrate topographic reconstruction, we acquired helium micrographs at five equally spaced azimuthal intervals of . Over-sampling of the system (equation (4)) was used to account for noisy data and/or for slight deviations from ideal diffuse scattering. In addition, the rotational stage of the sample manipulator has a small amount of mechanical inaccuracy, which limits the accuracy to which the center of rotation can be determined to a few pixels. Consequently, the tracking of the sample through rotations was imperfect and thus collecting more micrographs minimized the image area where fewer than three measurements were present, in total 24% of the image area could not be reconstructed.
The reconstruction process is illustrated in figure 4 and starts by manually identifying common points across the micrographs to correlate and align the separate azimuthal images. Distinct points on the sample were selected that are accurately identifiable across micrographs, as shown in panel 2. A transformation matrix, using only translational and rotational components, was calculated to create a mapping of each micrograph onto the same axis. Mapping to a common set of axes is crucial because the pixels in each image must correspond to the same point on the sample surface so that surface orientation may be inferred. Panel 3 shows the images after alignment, with 3 tracking points highlighted (out of a total 6 used for image alignment). Where at least 3 micrographs overlapped, the intensities from the micrographs were used to calculate the surface normals according to equation (4). Normal vectors for each pixel were generated, as shown in Figure 4 panel 4. Pixels that cannot be correlated across at least 3 micrographs are assigned a surface orientation parallel to the axis because otherwise the linear system, shown in equation (2), becomes underdetermined. Orientation parallel to the axis was chosen because it is the average of all possible orientations and thus underdetermined pixels have minimal effect on the integration of surface normal vectors which gives the final surface reconstruction. Finally, the normal map is integrated using the regularized least squares method[39, 40], giving the surface reconstruction, shown in panel 5. The reconstruction captures all qualitative aspects of the crystal: we are clearly observing the top half of a truncated octahedron.
Two metrics were employed to assess the quantitative accuracy of the reconstruction. First, the surface area of the top of the crystal was determined using 2D SHeM micrographs and independently via SEM (figure S1) to ensure the reconstruction process does not distort the size and shape of the sample. It should be noted that measurements taken directly from 2D SHeM micrographs and SEM images will both be affected by any tilt in the sample, or sample mounting. Second, the angles between the facets of the reconstructed surface were measured and compared to the known facet angles of an octahedral aluminum potassium sulfate crystal.
The dimensions measured for the size of the top face of the truncated pyramid are given in table 1. All of the dimensions agree to within experimental uncertainty, indicating that there is no significant distortion occurring due to the reconstruction algorithm.
| Dimensions of Pyramid Truncated Face | ||||
| SEM | ||||
| 2D SHeM | ||||
| 3D SHeM | ||||
The angle between the top facet of the crystal and the side facets for a truncated octahedron are (to 3 s.f.). To measure the angle between the reconstructed facets, the surface height going left to right across the crystal and top to bottom were averaged, producing two data sets each crossing the top of a trapezoid. Figure 5 shows a surface height map of the reconstruction in figure 4, with the extracted profiles horizontally and vertically across the crystal. Linear regression was applied to each edge of the two trapezoidal plots, the regression lines shown in colour in figure 5, resulting in four measurements of the facet angle; , , and , with a mean value of a overestimate compared to the ideal octahedron. The uncertainties in the linear regression were small, therefore the deviations from the expected value are not attributed to random noise.
Figures 4-5 reproduces the first demonstration of a 3D topographic map produced by neutral atoms, as presented by Radić et al.[25]. Moreover, our 3D reconstruction method can be applied to a single detector instrument, which is particularly significant as most SHeM instruments presented in the literature only utilize a single detector. However, the requirement for image alignment introduces an uncertainty in all dimensions of a few pixels, further limiting reconstruction accuracy. Due to the micrograph field-of-view being non-circular, when they are superimposed through rotation during the reconstruction process some pixels are lost as they become under-constrained and must be discounted. Using figure 4 as an example, 24% of the image area is unreconstructed, despite acquiring 5 micrographs. It should be possible to improve the ratio, and hence reduce the need for so many images, through more advanced scanning patterns and perhaps automated alignment during the acquisition process.
V Reconstruction using a multiple-detector instrument
Heliometric stereo on suitable multiple detector systems removes the need for rotational tracking entirely, but requires specialist hardware. The benefits of a multiple detector instrument are three-fold, 1) simultaneous imaging reduces imaging time by a factor of the number of detectors, 2) avoiding sample rotation increases reconstruction accuracy and further reduces imaging time because no sample alignment is required, and 3) decrease the likelihood that spatial drift in the sample or helium detection produces non-topographic contrast. We use the purpose-built multiple detector 2nd generation SHeM at Cambridge to perform our experiments (B-SHeM)[37]. The B-SHeM has the capacity for up to four helium detectors whose collection apertures can be arbitrarily placed around the sample owing to the versatility of the instrument’s defining optical element, the pinhole plate. The optical element can be 3D printed in plastic with complex internal geometries whilst retaining the (ultra)high-vacuum compatibility of traditionally machined metal using a simple baking method[42, 25].
In the presented work, we used the B-SHeM with three detectors (Hiden Analytical HAL/3F RC301 PIC30[43]) to minimally constrain the reconstruction and provide a lower-bound on the method’s accuracy. A schematic of the scattering geometry employed in the B-SHeM in the current work is shown in figure 6. Three elliptical detector apertures are placed with equiangular spacing around the scattering point, each forming a circular region of solid angle when viewed from the scattering point.
The reconstruction presented in this section marks a 6-fold decrease in acquisition time compared to a single detector implementation [25]. In the current work we also use a angle of incidence instead of normally incident beam in section IV. Non-normal incidence has significant implications for the reconstruction quality of vertical surfaces which will be discussed in detail in section VI.
To demonstrate multiple-detector heliometric stereo we selected a sample with known geometry for quantitative assessment of reconstruction quality. Two metal spheres with diameters were embedded in a substrate and imaged in the B-SHeM to obtain three simultaneous images that were then used to generate a surface profile using the heliometric stereo method. Figure 7 shows the acquired SHeM micrographs and the corresponding reconstructions of the spheres.
Using the known spherical geometry and dimension of the two spheres we can evaluate the quality of the heliometric stereo reconstruction via the overall root-mean-square (RMS) error, giving the total deviation of the reconstruction from the true shape. The RMS shape error is defined as the overall RMS error with the height difference contribution removed to provide a measure of the curvature of the sphere. Table 2 contains the calculated values of aspect ratios, overall RMS errors and RMS shape errors for spheres A and B, alongside the signal-to-noise ratios (SNR) for each of the three detectors.
The reconstruction of sphere A was qualitatively good with an overall RMS error calculated as 11.2% which exceeds the 3.4% predicted computationally by Lambrick et al.[26] for non-normal incidence scattering conditions. The difference can be attributed to two factors: (1) we use fewer detectors in the presented work than in the computational evaluation, 3 and 4, respectively. (2) The detectors used by B-SHeM have, on average, approximately a factor of 2 poorer SNR than the simulated detection where SNR=30 was set. Lambrick et al. found that the RMS error scales approximately with the SNR, which would then predict the 6. 8% RMS error with 4 images or the 7. 9% RMS error assuming the only impact of fewer images is a reduction in SNR (best case scenario), leaving 3.3% unaccounted for.
| Sample | Aspect Ratio |
| Det. 2 | Det. 3 | Det. 4 | Overall | Shape | ||||||
| Sphere A | 0.36 | 15.76 | 9.46 | 24.52 | 11.2% | 9.9% | ||||
| Sphere B | 0.52 | 17.20 | 14.37 | 33.28 | 13.8% | 13.9% | ||||
By contrast, the reconstruction for sphere B is qualitatively poorer. Indeed, Table 2 shows that the overall and shape errors are larger for sphere B. The greater size of sphere B means that it exhibits greater masking and shadowing in its SHeM micrographs and consequently decreasing reconstruction accuracy [26]. These effects are explored in more detail in the next section.
VI The effects of masking, shadowing and multiple scattering
The effects of masking, shadowing, and multiple scattering in SHeM micrographs all affect the quality of a heliometric stereo reconstructions. We define each of these terms as the following[20, 19, 24], with the mechanisms being shown diagrammatically in figure 8,
masking - an obstruction in the line-of-sight between the scattering site on the sample and the detector aperture,
shadowing - an obstruction in the line-of-sight between the incident beam aperture and a spot on the sample,
multiple scattering - any beam particle path that scatters from the sample more than once before detection.
As previously mentioned, these mechanisms break the key assumption of heliometric stereo: the direct relationship between scattered helium intensity and the surface orientation according to equation (1).
Masking, highlighted in area D of figure 7, causes dark regions in images, which will be reconstructed (equation (4)) as regions parallel to the detection direction. Shadowing results in areas of the sample where no information is gathered. Multiple scattering results in brighter regions of images, generally near vertical walls, recesses or similar features, such as in area C in figure 7, where a small recess at the edge of the embedded sphere results in enhanced contrast[24]. Bright regions caused by multiple scattering will also distort the solution to equation (4). the two spheres shown in figure 7 as our quantitative test of the impacts of masking and shadowing. The spheres provide a good test sample to quantify the effect of masking and shadowing on the method because both spheres have identical, and known, geometries, aside from sphere B having a larger aspect ratio because it is embedded in the substrate to a lesser degree. With masking and shadowing being geometrically defined as lines-of-sight between the sample and the detector or the incident aperture, respectively, when the aspect ratio of a feature increases the degree of masking and shadowing must also increase. This investigation is similar to that performed on simulated data in our previous work, where reconstruction accuracy was plotted as a function of the aspect ratio of features[26].
Figure 9 uses the known geometry of the spheres to plot the difference in reconstructed height, and the difference in reconstructed shape, from the expected geometry. Yellow regions lie higher than the expected surface, and conversely blue regions lie lower, green regions lie at the expected height.
For the lower aspect ratio sphere A, there is reasonably good accuracy across the sample, with only some small regions of poorer precision. In particular, the accuracy drops near the edge of the sphere, likely due to a crevice where the sphere is embedded into the surface, causing masking and multiple scattering. These errors are, however, locally confined, and do not cause errors away from the edges, similar to results found with simulated data[26]. Across the sphere itself the errors are asymmetric and match with the direction of the beam, but remain generally modest, generally reaching no more than double the average error.
We attribute the high quality reconstruction to the fact that the basic assumptions of heliometric stereo are held across almost the entire sample area: sphere A has a relatively low aspect ratio and no vertical faces, thus avoiding masking and exhibiting only relatively modest multiple scattering in the helium micrographs.
By contrast, while sphere B has the same basic shape as sphere A its increased aspect ratio results in more masking and shadowing. As can be seen in figure 9 the reconstructed surface lies noticeably higher towards the incident beam, and lower away from it, resulting in a visibly deformed reconstruction (figure 7). If we inspect quantitatively, we also find that the magnitude of the maximum deviation from the expected surface is also significantly greater than for sphere A, reaching and in percentage and shape, respectively. The greatest errors occur near the edge of the sphere, where there is non-negligible masking in all three micrographs, one of which is highlighted as feature D in figure 7, and where the surface of the sphere becomes approximately parallel to the beam.
Similar distortions were seen with simulated data when masking[26]; the simulated results suggest that by excluding masked regions from reconstruction more accurate shapes can be recovered, however, in order to exclude pixels and still perform a reconstruction more than 3 images are needed, which was not possible with the current experimental setup.
Sphere B also demonstrates the effects of multiple scattering. On the flat substrate near the sphere there is a slight distortion in the reconstruction, the surface is higher than expected, e.g. in highlighted region E in figure 9. Here, the intensity of the helium micrographs has been increased in a similar manner to that shown in figure 8 (c), causing distortions in the reconstruction.
As, in most circumstances, multiple scattering only accounts for a small fraction of SHeM contrast these distortions have remained modest. However, multiple scattering is intrinsic to neutral atom microscopy and can only be controlled to a limited degree with instrument geometry[19], thus future applications of heliometric-stereo will need to consider it carefully.
Shadowing, where the line-of-sight between incident beam and sample is blocked, is not present to a significant degree in the test samples considered here, however, it could lead to problems in reconstructions as parts of the sample could be missed. For a normal incidence configuration, such as that used in section IV, shadowing will occur at overhangs and vertical faces. For a non-normal incidence configuration, such as used in section V, vertical faces and overhangs can be captured if they face the beam, but regions facing away from the beam will be hidden. One approach to minimize the effect of shadowing would be to combining multiple reconstructions from different incidence conditions with the sample in different orientations, however, doing so would increase acquisition time.
Masking, which is intrinsic to neutral atom imaging, can be partially mitigated by changing the relative positions of the beam and the detectors: the larger the angular distance between the detector and beam, the greater the amount of masking[44]. An extreme optimization would be to place the detector in the same direction as the incident beam using an annular detector that surrounds the detector. The annular detection geometry is practicable experimentally with current polymer 3D printing methods used to create the pinhole plates for A-SHeM and B-SHeM. However, the annular detection geometry could only be used for one of the detectors so would either have to be combined with detectors that do exhibit masking.
All of these changes can, in principle, be implemented onto current single and multiple detector SHeM instruments cheaply and rapidly without significant hardware changes because the illumination and detection geometry is solely dictated by the pinhole plate optical element.
We include a reconstruction of a salt (sodium chloride) crystal in figure 10 using the same experimental setup as section V. Panels (a)-(c) show the helium micrographs used, panel (d) the resulting reconstruction and panel (e) an optical micrograph of the sample. The crystal has relatively low aspect ratio overall, but does posses a few vertical features at its edges that create strong shadowing or masking. Despite the local regions of shadowing and masking a reconstruction is generated that captures the overall topography of the sample well, highlighting the robustness of our method. By comparison to an optical microscope image (using a Keyence VHX-7000 microscope) we see that the heliometric stereo reconstruction reproduces the fine faceted structure on the surface of the crystal, including the cracks connecting the corners of the crystal, along with very fine ripples that appear parallel to the sides of the crystal forming concentric squares, which are of a size comparable to that of a single pixel in the SHeM micrographs. We note that these very fine features would not be recoverable with a single detector instrument as the lateral accuracy would be degraded by the image alignment procedure.
VII Conclusion
We present accurate surface profile reconstructions using both a single detector and a multiple detector instruments. Best results were obtained with samples suited to the technique: those that conform to the key assumptions of the method, most notably the assumption that the topography can be described by a single valued function . We expand upon the adaptation of the heliometric stereo method, as reported for a single detector instrument by Radić et al.[25], by presenting results to demonstrate the effect of excessive masking and shadowing due to samples with vertical edges and high aspect ratios. We suggest instrumentation changes to the illumination and detection geometry that can mitigate masking and shadowing, to improve reconstruction quality, that can be easily implemented into current SHeM instruments. We find, by comparing two different aspect ratio partial spheres, similarly to previous simulated results that reconstructions work well provided aspect ratios are kept .
Furthermore, we present a realistic sample of unknown geometry, the salt crystal in figure 10, and verify against optical microscopy that heliometric stereo successfully reproduces fine structures on the order of the pixel size. It was found that even when the assumptions were broken locally, e.g. by overhangs and vertical faces with the salt crystal sample, (features that prove challenging in many other techniques too) only local errors were introduced to the reconstructions.
Comparing the single and multiple detector reconstructions we highlight inherent advantages to the multiple detector implementation, including significantly faster imaging time, minimal manual intervention, and the recovering of pixel-level topographic information. With previously reported lateral SHeM spot sizes down to [45], and recent advances in neutral beam generation[46], existing SHeM instruments could, with minor modification, access sub-optical resolution surface reconstructions for moderate to high aspect ratio samples made of soft, insulating, optical transparent or electrically sensitive materials.
The success of the reconstruction process further validates recent work demonstrating that diffuse scattering is dominant for technological surfaces in SHeM. Looking ahead, we anticipate that the methods presented by Myles et al.[23] or the use of contrast features as presented by Lambrick et al.[24] will be combined with heliometric stereo for higher accuracy surface metrology, especially as SHeM resolution pushes into the nanoscale.
Supplementary information
VIII Single detector sample SEM image
IX Imaging parameters - single detector
The acquisition time for micrographs in figure 4 was 3h 24min per micrograph. Pixel sizes of were used with a dwell time of . pixels were used for the micrograph.
The acquisition time for the micrograph in figure 3 was 7h 16min. Pixel sizes of were used with a dwell time of . pixels were acquired – the version presented is cropped.
X Imaging parameters - multiple detector
The acquisition time for all three micrographs needed for a single reconstruction, as shown in figure 7 for two spheres, was approximately 8 hours for images with dimensions pixels and dwell time per pixel.
The acquisition time for all three micrographs needed for the reconstruction of a salt crystal in figure 10 was approximately 10.5 hours for images with dimensions pixels and dwell time per pixel.
Per surface reconstruction, the effective speed-up using the multiple detector instrument is compared to the single detector implementation, due to both simultaneous image acquisition and the removal of the manual sample alignment procedure.
Acknowledgements
The work was supported by EPSRC grant EP/R008272/1. The authors acknowledge support by the Cambridge Atom Scattering Centre (https://atomscattering.phy.cam.ac.uk) and EPSRC award EP/T00634X/1. The work was performed in part at corde, the Collaborative R&D Environment established to provide access to physics related facilities at the Cavendish Laboratory, University of Cambridge. SML acknowledges support from IAA award EP/X525686/1 and funding from Mathworks Ltd. We would like thank Boyao Liu for useful discussions. The authors acknowledge support from Ionoptika Ltd.
Author declarations
Conflict of interest
The authors have no conflicts of interest to report.
Data availability
Supporting data packs containing both raw and processed data, and Supplementary Material can be found at http://doi.org/10.17863/CAM.107537 for results with the Cambridge A-SHeM and https://doi.org/10.17863/CAM.114465 for results with the Cambridge B-SHeM.